For the reaction 2CH4(g)⇌C2H2(g)+3H2(g) 2 C H 4 ( g ) ⇌ C 2 H 2 ( g ) + 3 H 2 ( g ) K c = 0.135 at 1615 ∘C . What is Kp for the reaction at this temperature? Express your answer numerically.

For the reaction

N2(g)+3H2(g)⇌2NH3(g)N2(g)+3H2(g)⇌2NH3(g)

KpKp = 5.15×10−3 at 333 ∘C . What is Kc for the reaction at this temperature?

Enter your answer numerically.
Answer #1

2chy - Cl + 3 M Ke= 0.135 m T= 1615 + 273 = 1888 Kp = K (RT) Ang Kp = 0.135 (0.0821 X 1888) n = 4-2 =2 6135 X Kp = 3243.5 [kp